# Revisiting Euclidean Alignment for Transfer Learning in EEG-Based Brain-Computer Interfaces

### v1

@@ 2.3 Applications of EA
In the original paper [13], the effectiveness of EA was only demonstrated in two classic BCI paradigms, motor imagery and P300 visual evoked potentials, using traditional classifiers. In the past five years, researchers from all over the world have applied it to eight additional BCI paradigms, including mental imagery, rapid series visual presentation (RSVP), error-related negativity, sleep stage classification, stress detection, emotion recognition, motor execution, and seizure detection, using both traditional classifiers and deep neural networks, all achieving performance improvements, as summarized in Table 1. These 10 BCI paradigms cover almost all non-invasive BCI applications, i.e., EA has very broad applicability.
Table 1: Performance improved by EA in 10 different BCI paradigms.
BCI Paradigm Classifier Accuracy (%) Reference
without EA with EA Improvement
Motor Imagery CSP-LDA 59.71 79.79 20.08 He and Wu, IEEE TBME 2020 [13]
@@
Mental Imagery MDM 56.02 66.08 10.06 Kumar et al., BCI 2019 [51]
Event-Related Potential xDAWN-SVM 64.60 68.80 4.20 He and Wu, IEEE TBME 2020 [13]
RSVP CSP-LDA 65.36 69.07 3.71 Zhang and Wu, IEEE TNSRE 2020 [22]
Error-Related Negativity CSP-LDA 61.87 64.63 2.76
Sleep Stage Classification EEGInception 56.87 67.70 10.83 Bakas et al., NeurIPS 2021 [23]
Stress Detection SDCAN 45.69 48.17 2.48 Fu et al., IEEE TNSRE 2022 [52]
Emotion Recognition AM-MSFFN 88.00 93.51 5.51 Jiang et al., IEEE SJ 2023 [53]
SSVEP SUTL 72.28 74.04 1.76 Li et al., ESWA 2024 [54]
Motor Execution EEGNet 52.10 62.50 10.40 Bakas et al., JNE 2025 [40]
Seizure Detection EEGNet 68.43 75.86 7.43 Wang et al., NSR 2025 [55]
@@
3 Extension of EA
This section introduces LA, an extension of the unsupervised EA to supervised EEG data alignment.
LA [17] was also proposed by the author’s group in 2020, primarily for EEG-based motor imagery classification. LA assumes all source domain samples are labeled, and the target domain also has a small number of labeled calibration samples. The label information could be properly used for better EEG trial alignment.
The idea of LA is shown in Fig. 1. It essentially performs class-wise EA.
For simplicity, assume the source domain and the target domain have the same label space, e.g., for motor imagery, the source domain has two classes (left hand and right hand), and the target domain has the same two classes. For a given class, assume the source domain has $N_{\mathrm{s}}$ labeled EEG trials $\{(X_{n}^{\mathrm{s}},y_{n}^{\mathrm{s}})\}_{n=1}^{N_{\mathrm{s}}}$, and the target domain has $N_{\mathrm{t}}$ labeled EEG trials $\{(X_{n}^{\mathrm{t}},y_{n}^{\mathrm{t}})\}_{n=1}^{N_{\mathrm{t}}}$.
@@ 3 Extension of EA
After LA, the class-wise mean covariance matrix of all $N_{\mathrm{s}}$ aligned source domain trials $\{\tilde{X}_{n}^{\mathrm{s}}\}_{n=1}^{N_{\mathrm{s}}}$ equals the class-wise mean covariance matrix of the $N_{\mathrm{t}}$ target domain trials $\{\tilde{X}_{n}^{\mathrm{t}}\}_{n=1}^{N_{\mathrm{t}}}$:
@@
i.e., for each class, the mean covariance matrices of both domain are equal, and hence the class-wise EEG data distributions from different domains become more consistent.
@@
LA performs the above alignment for each class in the source domain separately. Note that it transforms the source domain trials to match the target domain trials. Unlike in EA that all EEG trials in the same domain share the same reference matrix, in LA different classes in the same domain use different reference matrices. When LA is used in testing, we do not know the class label (this is to be predicted) for an incoming EEG trial, so we cannot determine which reference matrix to use. By transforming the known source domain trials only, we can avoid this problem.
@@
We [17] showed that LA can also be used when the source and target domains have different label spaces. For example, when the source domain has two classes of left hand and right hand, and the target domain has two classes of feet and tongue, we can align the source domain left hand class with the target domain feet class, and the source domain right hand class with the target domain tongue class, still achieving effective TL. Of course, EA could also be used in this challenging scenario, as it does not distinguish among the classes at all.
Generally, LA outperforms EA, as it exploits the useful label information to make aligned trials from different classes more separable.
@@
LA has also demonstrated its effectiveness in multiple BCI paradigms, including motor imagery classification [17, 15, 57, 58], emotion recognition [59, 60, 61], mental workload classification [62], speech imagery classification [63], and traumatic brain injury classification [64].
4 New Research Directions
@@
4.1 Improvements to EA
Though EA has achieved promising performance in 10 different BCI paradigms, it still has some limitations, which could be address to further improve its performance:
  1. 1.
    EA needs to compute the inverse of a matrix, i.e., $\bar{R}^{-1/2}$ in (2), which may be unstable when the number of channels is large.
  2. 2.
    The original EEG trials from different subjects have consistent channel configurations before alignment, because they use the same EEG headset. For example, for the MI trials in our experiments, the first channel was always AF3, the second channel was AF4, and the 59th channel was O2. However, after EA, the aligned channels are no longer consistent, e.g., the first aligned channel of Subject 1 may be obtained by applying weights $[0.1,0.2,...,0.5]$ to the 59 original channels, whereas the first aligned channel of Subject 2 may be obtained by applying weights $[0.3,0.1,...,0.2]$ to the 59 original channels. This channel location shifting may impact the subsequent classification performance.
4.2 Applicability of EA to Other Multi-Channel Physiological Signals
@@
Since many other physiological signals, e.g., Magnetoencephalography (MEG) [67], Electrocardiography (ECG) [68] and stereo-Electroencephalography (sEEG) [69], also have multiple channels and large individual differences, it is interesting to study if EA can also be applied to them to reduce the data distribution discrepancies among subjects.
4.3 Large Models for EEG Decoding
@@
4.4 Extension from Classification to Regression
So far, the applications of EA exclusively focused on BCI classification problems, such as motor imagery, event related potential, seizure classification, etc., as shown in Table 1. However, there are also many important regression problems in BCIs [11], such as primitive emotion estimation [4], driver drowsiness estimation [72], user reaction time estimation [73], etc. As in BCI classification problems, EEG signal stationarity and large individual differences also exist in BCI regression problems. So, TL could also be used to facilitate the calibration [74].
Our previous research has extended some blocks in the TL pipeline for motor imagery classification (Fig. 2), e.g., common spatial pattern filtering [75], wAR classifier [74], and more traditional and deep TL classifiers [76], to BCI regression problems, using fuzzy sets. It is interesting to study if EA, and also LA, could be extended from classification to regression, also using fuzzy sets.
4.5 Accurate and Robust EEG Decoding
@@
Our recently proposed alignment based adversarial training (ABAT) algorithm [87] solved this challenging problem. Its flowchart is shown in Fig. 5. Essentially, it first uses EA to reduce the data distribution discrepancies among different domains, and hence to improve the classification accuracy on benign EEG trials; it then performs adversarial training to improve the robustness on adversarial samples.
Figure 5: Alignment based adversarial training (ABAT).
@@
In summary, EA is essential in adversarial training to simultaneously improve the decoding accuracy and adversarial robustness of EEG-based BCIs. It is expected that more sophisticated and comprehensive TL pipeline, like the one shown in Fig. 2, can be integrated with adversarial training to achieve better performance.
4.6 Accurate, Robust, and Privacy-Preserving EEG Decoding
@@
Various approaches, including security multiparty computation [91], synthetic data generation [92, 93], homomorphic encryption [94], source-free TL [95, 96, 97, 98], federated learning [99], machine unlearning [100], and data perturbation [101], have been proposed for privacy-preserving BCIs.
@@
The flowchart of using A3E in EEG-based motor imagery classification is shown in Fig. 6. Each source domain uses EA to align their EEG trials first. Since EA uses information from its own domain only, there is no privacy issues. Three privacy protection scenarios are then considered:
@@
5 Conclusions
Due to the non-stationarity and large individual differences of EEG signals, EEG-based BCIs usually need subject-specific calibration to tailor the decoding algorithm for each new subject, which is time-consuming and user-unfriendly, hindering their real-world applications. TL has been extensively used to expedite the calibration, by making use of EEG data from other subjects/sessions. An important consideration in TL for EEG-based BCIs is to reduce the data distribution discrepancies among different subjects/session, to avoid negative transfer. EA was proposed in 2020 to address this challenge. Numerous experiments from 10 different BCI paradigms demonstrated its effectiveness and efficiency. This paper has revisited the EA, explaining its procedure and correct usage, introducing its applications and extensions, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.

### v2

@@ II-E Applications of EA
In the original paper [13], the effectiveness of EA was only demonstrated in two classic BCI paradigms, motor imagery and P300 visual evoked potentials, using traditional classifiers. In the past five years, researchers from all over the world have applied it to 10 additional BCI paradigms, including mental imagery, rapid series visual presentation (RSVP), error-related negativity, sleep stage classification, motor+speech imagery, stress detection, emotion recognition, steady-state visual evoked potential (SSVEP), motor execution, and seizure detection, using both traditional classifiers and deep neural networks, all achieving promising performance improvements, as summarized in Table III. These 13 BCI paradigms cover almost all non-invasive BCI applications, i.e., EA has very broad applicability.
TABLE III: Classification accuracy improved by Euclidean alignment in 13 different BCI paradigms. Note that different publications generally used different datasets.
No. BCI Paradigm Classifier Accuracy (%) Reference
w/o EA w/ EA Improvement
1 Motor Imagery CSP-LDA 59.71 79.79 20.08 He and Wu, IEEE TBME 2020 [13]
@@
BaseNet 57.90 71.00 13.20 Wimpff et al., arXiv [66]
2 Mental Imagery MDM 56.02 66.08 10.06 Kumar et al., BCI 2019 [18]
3 Event-Related Potential xDAWN-SVM 64.60 68.80 4.20 He and Wu, IEEE TBME 2020 [13]
4 RSVP CSP-LDA 65.36 69.07 3.71 Zhang and Wu, IEEE TNSRE 2020 [42]
5 Error-Related Negativity CSP-LDA 61.87 64.63 2.76
6 Sleep Stage Classification EEGInception 56.87 67.70 10.83 Bakas et al., NeurIPS 2021 [43]
7 Motor+Speech Imagery DS-CSP 63.62 67.09 3.47 Zhan et al., CBM 2022 [34]
DS-TSM 66.12 68.42 2.30
8 Stress Detection SDCAN 45.69 48.17 2.48 Fu et al., IEEE TNSRE 2022 [67]
9 Emotion Recognition AM-MSFFN 88.00 93.51 5.51 Jiang et al., IEEE SJ 2023 [68]
10 SSVEP SUTL 72.28 74.04 1.76 Li et al., ESWA 2024 [69]
11 Motor Execution EEGNet 52.10 62.50 10.40 Bakas et al., JNE 2025 [55]
12 Seizure Detection EEGNet 68.43 75.86 7.43 Wang et al., NSR 2025 [70]
13
Traumatic Brain
Injury Classification
CNN 61.00 89.00 28.00 Vishwanath et al., BSPC 2025 [71]
@@
EA has also achieved remarkable performance in BCI competitions.
By integrating it into our TL algorithms, the author’s team won nine National Championships in China BCI Competition (the most prestigious BCI competition in China, organized jointly by the National Natural Science Foundation of China, Chinese Institute of Electronics, and Tsinghua University) between 2019 and 2024.
In the Benchmarks for EEG Transfer Learning [^7] https://beetl.ai/challenge (BEETL) competition organized at NeurIPS 2021, the champion team’s approach [72]for improving subject-independence is to align the latent feature distributions of the deep learning models between different subjects and sessions. This is similar to the Euclidean Alignment method, which performs spatial whitening with the average covariance matrix per subject, although the methods are not constrained to be applied on the model input.” In the runner-up’s approach [72], “Euclidean Alignment is used to close the gap between subjects in source domains and target domain. All trials of each subject are aligned such that the mean covariance matrices of all subjects are equal to the identity matrix after alignment. After alignment, data set distribution in both source domains and target domain are more similar.
III Extension/Counterparts of EA
This section introduces LA, an extension of the unsupervised EA to supervised EEG data alignment, and several other EEG data alignment approaches.
III-A Label Alignment
LA [15] was also proposed by the author’s group in 2020, primarily for heterogeneous EEG-based motor imagery classification, i.e., the motor imagery classes for different subjects are partially or completely different. LA assumes all source domain samples are labeled, and the target domain also has a small number of labeled calibration samples. The label information could be properly used for better EEG trial alignment.
The idea of LA is shown in Figure 1. It essentially performs class-wise EA to reduce the class conditional EEG data distribution discrepancies between the source and target domains.
As an example, assume the source subject imagines the movement of left hand or right hand, whereas the target subject imagines the movement of left hand or both feet. Then, we pair the source subject left hand motor imagery movement with the target subject left hand movement, and the source subject right hand movement with the target subject both feet movement. For a given class-pair, assume the source domain has $N_{\mathrm{s}}$ labeled EEG trials $\{(X_{n}^{\mathrm{s}},y_{n}^{\mathrm{s}})\}_{n=1}^{N_{\mathrm{s}}}$, and the target domain has $N_{\mathrm{t}}$ labeled EEG trials $\{(X_{n}^{\mathrm{t}},y_{n}^{\mathrm{t}})\}_{n=1}^{N_{\mathrm{t}}}$.
@@ III-A Label Alignment
Here $\bar{R}_{\mathrm{t}}^{1/2}\bar{R}_{\mathrm{s}}^{-1/2}$ is called the transformation matrix in LA.
After LA, the class-wise mean covariance matrix of all $N_{\mathrm{s}}$ aligned source domain trials $\{\tilde{X}_{n}^{\mathrm{s}}\}_{n=1}^{N_{\mathrm{s}}}$ equals the paired class-wise mean covariance matrix of the $N_{\mathrm{t}}$ target domain trials $\{\tilde{X}_{n}^{\mathrm{t}}\}_{n=1}^{N_{\mathrm{t}}}$:
@@
i.e., for each paired class, the mean covariance matrices of both domain are equal, and hence the class conditional EEG data distributions from different domains become more consistent.
@@
LA performs the above alignment for each paired class in the source domain separately. Note that it transforms the source domain trials to match the target domain trials. Unlike in EA that all EEG trials in the same domain share the same transformation matrix, in LA different classes in the same domain use different transformation matrices. When LA is used in testing, we do not know the class label (this is to be predicted) for an incoming target domain EEG trial, so we cannot determine which transformation matrix to use. By transforming the known source domain trials only, we can avoid this problem.
@@
We [15] showed that LA can also be used when the source and target domains have completely different label spaces. For example, when the source domain has two classes of left hand and right hand, and the target domain has two classes of feet and tongue, we can align the source domain left hand class with the target domain feet class, and the source domain right hand class with the target domain tongue class, still achieving effective TL. Of course, EA could also be used in this challenging scenario, as it does not distinguish among the classes at all.
@@
LA has demonstrated promising performance in multiple BCI paradigms, particularly in heterogeneous TL, including motor imagery classification [15, 72, 73, 74], motor+speech imagery classification [34], traumatic brain injury classification [75], emotion recognition [76, 77, 78], and mental workload classification [79].
TABLE IV: Classification accuracy improved by label alignment in 5 different BCI paradigms. Note that different publications generally used different datasets.
No. BCI Paradigm Classifier Accuracy (%) Reference
w/o LA w/ LA Improvement
1 Motor Imagery CSP-LDA 51.10 69.95 18.85 He and Wu, IEEE TNSRE 2020 [15]
TS-SVM 56.65 71.75 15.10
2 Motor+Speech Imagery DS-CSP 63.62 66.64 3.02 Zhan et al., CBM 2022 [34]
DS-TSM 66.12 71.12 5.00
3 Traumatic Brain Injury Classification DT 62.24 65.78 3.54 Vishwanath et al., IEEE EMBC 2022 [75]
RF 67.93 70.18 2.25
kNN 71.31 74.42 3.11
MLP 73.66 78.73 5.07
SVM 74.46 74.86 0.40
XGB 66.89 69.41 2.52
4 Emotion Recognition TCA 46.72 68.08 21.36 Peng et al., IEEE TII 2023 [76]
kNN 48.14 66.46 18.32 Ren et al., KBS 2024 [77]
GFK 42.80 68.10 25.30 Zhang et al., IEEE JBHI 2024 [78]
5
Mental
Workload
Classification
LDA 56.63 82.53 25.90 Wang et al., IEEE JBHI 2024 [79]
LA may outperform EA when the source and target domains have different label spaces. For example, we [15] demonstrated that when the source subject imagines the movements of right hands or both feet, whereas the target subject imagines the movements of left hand or tongue, the baseline CSP-LDA (linear discriminant analysis) classification accuracy without any alignment was 51.10%, the classification accuracy of EA-CSP-LDA was 64.01%, and the classification accuracy of LA-CSP-LDA was 69.95%.
However, when the source and target domains have the same label space, EA may outperform LA. For example, Figure 7 shows the performance comparison of TF-CAR-EA-RCSP-wAR (the best-performer in Figure 4) and TF-CAR-LA-RCSP-wAR, which replaces EA by LA. LA performed much worse than EA, because given a small number of labeled target domain samples, it is difficult to accurately estimate $\bar{R}_{\mathrm{t}}^{1/2}$ in the target domain, and hence the transformation matrix $\bar{R}_{\mathrm{t}}^{1/2}\bar{R}_{\mathrm{s}}^{-1/2}$. He and Wu [15] used $k$-medoids clustering to select $k$ most representative unlabeled EEG trials to label, making the estimation of $\bar{R}_{\mathrm{t}}^{1/2}$ more reliable than random selection, improving the performance of LA.
In summary, when the source and target domains have the same label space, EA is preferred over LA; otherwise, LA is preferred over EA.
Figure 7: The motor imagery classification performance of six different transfer learning pipelines. TF: Temporal filtering; RCSP: Regularized common spatial pattern; wAR: weighted adaptation regularization; CAR: Common average referencing; EA: Euclidean alignment; LA: Label alignment.
III-B Comparison of EEG Data Alignment Approaches
In addition to Riemannian alignment, EA and LA, there are several other EEG data alignment approaches proposed in the literature, which can be categorized into two classes:
  1. 1.
    Alignments of the EEG trials by transformation matrices, and the aligned EEG trials can then be used in both traditional machine learning and deep learning. Representative approaches include Riemannian Alignment [16], Parallel Transport [80], Riemannian Procrustes Analysis [33], EA [13], Centroid Alignment [42], and LA [15].
  2. 2.
    Alignments of the mini-batches in deep learning, which are specific to deep learning. Representative approaches include Adaptive Batch Normalization (adaBN) [81, 82] and Latent Alignment [55].
Table I in our previous publication [11] gives a comprehensive comparison of the first category of EEG data alignment approaches. For the completeness of this paper, we collect and update its results, and add adaBN and latent alignment, as shown in Table V.
TABLE V: Comparison of different EEG data alignment approaches [11].
Criterion
Riemannian
Alignment
[16]
Parallel
Transport
[80]
Riemannian
Procrustes
Analysis [33]
Euclidean
Alignment
[13]
Centroid
Alignment
[42]
Label
Alignment
[15]
adaBN
[81, 82]
Latent
Alignment
[55]
Applicable
Paradigms
7 different
paradigms
(Table I)
MI [80]
MI [33],
ERP [33],
SSVEP [33]
13 different
paradigms
(Table III)
MI [42],
ERP [42],
Emotion
recognition
[76, 77, 78]
5 different
paradigms
(Table IV)
MI [82],
Cognitive load
recognition
[81]
MI [55],
ERP [55],
Sleep stage
classifica-
tion [55]
Online or
Offline
Both Both Both Both Both Offline Both Both
Need
Labeled Target
Domain Trials
No for MI,
Yes for others
No Yes No No Yes No No
What to
Align
Riemannian
space
covariance
matrices
Riemannian
Tangent
space
features
Riemannian
space
covariance
matrices
Euclidean
space
EEG
trials
Riemannian
space
covariance
matrices
Euclidean
space
EEG
trials
Euclidean
space EEG
features
in a
mini-batch
Euclidean
space EEG
features
in a
mini-batch
Reference
Matrix
Calculation
Riemannian
mean of
resting or
non-target
trial
covariance
matrices in
each domain
Riemannian
mean of all
covariance
matrices in
each domain
Riemannian
mean of all
labeled
covariance
matrices in
each domain
Euclidean
mean of all
covariance
matrices in
each domain
Riemannian,
Euclidean, or
Log-Euclidean
mean of all
covariance
matrices in
each domain
Log-Euclidean
mean of
labeled
covariance
matrices
in each class
of each domain
Unneeded Unneeded
Classifier
Riemannian
space only
Euclidean
space only
Riemannian
space only
Riemannian or
Euclidean space
Riemannian or
Euclidean space
Riemannian or
Euclidean space
Euclidean
space only
Euclidean
space only
Handle Class
Mismatch
between
Domains
No No No No No Yes No No
Computational
Cost
High High High Low Low Low Low Low
In application domains other than BCIs, many classic data alignment approaches for TL have also been proposed [83], e.g., transfer component analysis (TCA) [84] and CORrelation ALignment (CORAL) [85].
TCA assumes that both marginal and conditional probability distributions of the source domain $X_{\mathrm{s}}$ and target domain $X_{\mathrm{t}}$ are different, and tries to find a transformation matrix $W$ to make them approximately identical. More specifically, $W$ needs to satisfy two constraints [84]: 1) the distance between the marginal distributions of $WX_{\mathrm{s}}$ and $WX_{\mathrm{t}}$ is small; and, 2) $WX_{\mathrm{s}}$ and $WX_{\mathrm{t}}$ preserve important properties of $X_{\mathrm{s}}$ and $X_{\mathrm{t}}$. Compared with TCA, EA only considers the first constraint but not the second, which could be a direction for further improvement.
CORAL minimizes the domain shift by aligning the second-order statistics of their distributions. Assume a $d$-dimensional feature vector has been extracted for each sample in each domain. CORAL first computes the feature covariance matrices $C_{\mathrm{s}}\in\mathbb{R}^{d\times d}$ and $C_{\mathrm{t}}\in\mathbb{R}^{d\times d}$ in the source domain and target domain, and then finds a transformation matrix $W\in\mathbb{R}^{d\times d}$ for the source domain such that the Frobenius norm of the difference between the covariance matrices, $||W^{\top}C_{\mathrm{s}}W-C_{\mathrm{t}}||_{F}^{2}$, is minimized. As pointed out in [13, 15], the main difference between CORAL and EA/LA is that: “CORAL uses 1D features, and each domain has only one feature covariance matrix, which measures the covariances between different pairs of individual features. EA/LA uses 2D features (EEG trials), and each EEG trial has a covariance matrix, which measures the covariances between different pairs of EEG channels. So, the covariance matrices in CORAL and EA/LA have different meanings.” Meanwhile, CORAL may also be used to align the log-variance features in TF-CAR-EA-RCSP-wAR before they are fed into wAR for classification.
IV New Research Directions
@@
IV-A Applicability of EA to Other Multi-Channel Physiological Signals
@@
Since many other physiological signals, e.g., Magnetoencephalography (MEG), Electrocardiography (ECG) and stereo-Electroencephalography (sEEG), also have multiple channels and large individual differences, it is interesting to study if EA can also be applied to them to reduce the data distribution discrepancies among subjects.
IV-B Large Models for EEG Decoding
@@
IV-C Extension from Classification to Regression
So far, the applications of EA exclusively focused on BCI classification problems, such as motor imagery, event related potential, seizure classification, etc., as shown in Table III. However, there are also many important regression problems in BCIs [11], such as primitive emotion estimation [4], driver drowsiness estimation [90], user reaction time estimation [91], etc. As in BCI classification problems, EEG signal stationarity and large individual differences also exist in BCI regression problems. So, TL could also be used to facilitate the calibration [92].
Our previous research has extended some blocks in the TL pipeline for motor imagery classification (Figure 2), e.g., common spatial pattern filtering [93] and wAR classifier [92], to BCI regression problems, using fuzzy sets. It is interesting to study if EA/LA could be extended from classification to regression, also using fuzzy sets.
IV-D Accurate and Robust EEG Decoding
@@
Our recently proposed alignment based adversarial training (ABAT) algorithm [104] solved this challenging problem. Its flowchart is shown in Figure 8. Essentially, it first uses EA to reduce the data distribution discrepancies among different domains, and hence to improve the classification accuracy on benign EEG trials; it then performs adversarial training to improve the robustness on adversarial samples.
Figure 8: Alignment based adversarial training (ABAT) to simultaneously increase the decoding accuracy and adversarial robustness.
@@
In summary, EA is essential in adversarial training to simultaneously improve the decoding accuracy and adversarial robustness of EEG-based BCIs. It is expected that more sophisticated and comprehensive TL pipeline, like the one shown in Figure 2, can be integrated with adversarial training to achieve better performance.
IV-E Accurate, Robust, and Privacy-Preserving EEG Decoding
@@
Various approaches, including security multiparty computation [108], synthetic data generation [109, 110], homomorphic encryption [111], source-free TL [26, 112, 113, 114], federated learning [115], machine unlearning [116], and data perturbation [117, 118], have been proposed for privacy-preserving BCIs.
@@
The flowchart of using A3E in EEG-based motor imagery classification is shown in Figure 9. Each source domain uses EA to align their EEG trials first. Since EA uses information from its own domain only, there is no privacy issues. Three privacy protection scenarios are then considered:
@@
V Conclusions
Due to the non-stationarity and large individual differences of EEG signals, EEG-based BCIs usually need subject-specific calibration to tailor the decoding algorithm for each new subject, which is time-consuming and user-unfriendly, hindering their real-world applications. TL has been extensively used to expedite the calibration, by making use of EEG data from other subjects/sessions. An important consideration in TL for EEG-based BCIs is to reduce the data distribution discrepancies among different subjects/session, to avoid negative transfer. EA was proposed in 2020 to address this challenge. Numerous experiments from 13 different BCI paradigms demonstrated its effectiveness and efficiency. This paper has revisited EA, explaining its procedure and correct usage, introducing its applications and extensions, and pointing out potential new research directions. It should be very helpful to BCI researchers, especially those who are working on EEG signal decoding.